# SymTFT out of equilibrium: from time crystals to braided drives and Floquet codes

### 28 Dec 2023 v1

@@ Appendix A Hilbert space, charges, boundary conditions, degeneracies
In the main text we have discussed how the Hilbert space $\mathcal{H}_{\text{dyn}}$ of the 1D system is embedded into the 2D topological order (TO) as the subspace with no anyons away from the boundary $B_{\text{dyn}}$. Here we discuss how $\mathcal{H}_{\text{dyn}}$ decomposes into symmetry sectors and is enlarged with twisted sectors. We also provide further discussion on twisted boundary conditions (BCs) and spectral degeneracies. Our approach bridges between the condensed-matter oriented Refs. Lichtman et al., 2021; Moradi et al., 2023, with Ref. Moradi et al., 2023 working on the infinite cylinder and considering twisted BCs, while Ref. Lichtman et al., 2021 working on a finite cylinder (thus having SymTFT access to certain spectral degeneracies) but without twisted BCs. Ref. Gaiotto and Kulp, 2021 establishes results analogous to those discussed in this Appendix from a high-energy theory perspective. See also Refs. Chatterjee and Wen, 2023; Ji and Wen, 2020; Bhardwaj and Schafer-Nameki, ; Bhardwaj et al., ; Bhardwaj et al., .
@@
The Hilbert space, for finite Abelian $G$, is $\mathcal{H}_{\text{dyn}}=\mathcal{H}_{1}^{\otimes L}$, with $\mathcal{H}_{1}=\text{span}(\ket{g},g\in G)$. It is decomposed into symmetry sectors as $\mathcal{H}_{\text{dyn}}=\oplus_{\alpha}\mathcal{H}_{\text{dyn}}^{(1,\alpha)}$, where subspaces $\mathcal{H}_{\text{dyn}}^{(1,\alpha)}$ are of states “charged” according to $\alpha\in\text{Rep}\,G$ under the $G$ symmetry, $\smash{\overline{W}}^{(g,1)}\ket{\psi}=\chi_{\alpha}(g)\ket{\psi}$ for all |$\ket{\psi}\in\mathcal{H}_{\text{dyn}}^{(1,\alpha)}$. Starting with the $G$-symmetric subspace $\mathcal{H}_{\text{dyn}}^{(1,1)}$, satisfying $\smash{\overline{W}}^{(g,1)}\ket{\psi}=\ket{\psi}$, $\forall g\in G,\forall\ket{\psi}\in\mathcal{H}_{\text{dyn}}^{(1,1)}$, by $\smash{\overline{W}}^{(g,1)}W_{k}^{(1,\alpha)}=\chi_{\alpha}(g)W_{k}^{(1,\alpha)}\smash{\overline{W}}^{(g,1)}$, we have $\mathcal{H}_{\text{dyn}}^{(1,\alpha)}=W_{k}^{(1,\alpha)}\mathcal{H}_{\text{dyn}}^{(1,1)}$ (regardless of the choice $k$ of endpoint). The operators $W_{k}^{(1,\alpha)}$ thus introduce charge $(1,\alpha)$ to $B_{\text{dyn}}$ and thus to the 1D system.
@@
The space $\mathcal{H}_{\text{dyn}}^{(g,\alpha)}=W_{\tilde{j}}^{(g,1)}\mathcal{H}_{\text{dyn}}^{(1,\alpha)}=W_{\tilde{j}}^{(g,1)}W_{k}^{(1,\alpha)}\mathcal{H}_{\text{dyn}}^{(1,1)}$ is of states obeying $g$-twisted BCs. In the main text we interpreted this via the “$G$-Aharonov Bohm” effect; here we provide a more direct interpretation: Consider $\mathcal{T}_{L}$, the translation of $B_{\text{dyn}}$, and with it $\tilde{j}$ and $k$, by $L$ sites, keeping the rest of the cylinder fixed. This wraps $W_{\tilde{j}}^{(g,1)}W_{k}^{(1,\alpha)}\equiv W^{(a)}$ [with $a=(g,\alpha)$] around the cylinder and, since reconnecting the resulting $a$ string (more precisely, ribbon) amounts to a braid, we have $\mathcal{T}_{L}W^{(a)}\mathcal{T}_{L}^{\dagger}=e^{i\theta_{aa}}W^{(a)}\smash{\overline{W}}^{(a)}$. Thus, by $\smash{\overline{W}}^{(a)}\ket{\psi}=\mathcal{T}_{L}^{\dagger}\ket{\psi}=\ket{\psi}$ for |$\ket{\psi}\in\mathcal{H}_{\text{dyn}}^{(1,1)}$, we find that $\mathcal{T}_{L}\ket{\psi}=\chi_{\alpha}(g)\ket{\psi}$ for |$\ket{\psi}\in\mathcal{H}_{\text{dyn}}^{(g,\alpha)}$: the BCs are twisted. Note that the link $\tilde{j}$ where we insert the BC twist is a gauge choice: it can be moved to link $\tilde{j}^{\prime}$ upon conjugating the Hamiltonian (or Floquet unitary) by $W_{jj^{\prime}}^{(g,1)}$. To consider all twisted sectors together, we may work in the enlarged Hilbert space ¯$\overline{\mathcal{H}}_{\text{dyn}}=\oplus_{\alpha,g}\mathcal{H}_{\text{dyn}}^{(g,\alpha)}$. When different BCs are coupled, as is the case in the presence of $\tilde{G}$-breaking perturbations, we must work with ¯$\overline{\mathcal{H}}_{\text{dyn}}$.
@@
Turning to degeneracies, first recall that a gapped, or more generally many-body localized (MBL), 1D system is characterized by a Lagrangian subgroup $\mathcal{A}$; the $W_{k,k+1}^{(a)}$ with $a\in\mathcal{A}$ form a maximal set of local, mutually commuting $G$-symmetric operators, defining the zero-correlation-length (or fixed-point, FP) limit of the gapped, or MBL, phase. The topological ground-state degeneracy on the cylinder, equal to $|\mathcal{A}\cap\mathcal{A}_{\text{ref}}|$, i.e., the number of distinct strings that can connect $B_{\text{ref}}$ and $B_{\text{dyn}}$ without creating anyons, while avoiding crossing domain wall $D^{(\sigma)}_{\text{ref}}$, yields the ground-state (or spectral, for MBL) degeneracy of the 1D system within a twisted sector $\oplus_{\alpha}\mathcal{H}_{\text{dyn}}^{(g,\alpha)}$. Strings $W_{\tilde{j}}^{(g,1)}$ that can similarly connect $B_{\text{ref}}$ and $B_{\text{dyn}}$ while crossing $D^{(\sigma)}_{\text{ref}}$ yield degeneracies between twisted sectors. In the Floquet setting, both types of degeneracies may split into topological spectral multiplets (such as $\pi$ spectral pairing as discussed for $G=\mathbb{Z}_{2}$), with structure set by the drive’s $\smash{\overline{W}}^{(b)}$.
@@ Appendix B Absolute stability in the $0\pi$PM
To study this, we consider $U_{\text{F}}^{(e,\langle m\rangle)}=\smash{\overline{W}}^{(e)}e^{-iH_{1}}e^{-iH_{0}}$ with $H_{0}=\sum_{j}J_{j}X_{j}+J_{j}^{(\text{int})}X_{j}X_{j+1}+\tilde{h}\tilde{X}_{\tilde{1}}$ and $H_{1}=\sum_{j}h_{j}W_{j,j+1}^{(e)}+h_{j}^{(\text{int})}W_{j,j+2}^{(e)}$. With only $\tilde{X}_{\tilde{1}}$ as link operator in $H_{0}$, it is sufficient to just gauge link $\tilde{1}$, i.e. to take $W_{j,j+1}^{(e)}=Z_{j}Z_{j+1}$ for $j\neq 1$ and $W_{1,2}^{(e)}=Z_{1}\tilde{Z}_{\tilde{1}}Z_{2}$, and $W^{(e)}_{L}=\tilde{Z}_{\tilde{1}}$. Since $\smash{\overline{W}}^{(e)}\tilde{X}_{\tilde{1}}=-\tilde{X}_{\tilde{1}}\smash{\overline{W}}^{(e)}$, the $\tilde{h}\tilde{X}_{\tilde{1}}$ term breaks ~$\tilde{\mathbb{Z}}_{2}$ symmetry; it couples the system’s untwisted, i.e., periodic boundary condition (PBC), and $m$-twisted, i.e., anti-periodic boundary condition (APBC), sectors.
@@
We take $J_{j}$ distributed uniformly within $[\bar{J}/2,3\bar{J}/2]$ and $h_{j}$ within $[-\lambda\bar{J},\lambda\bar{J}]$ (with $\lambda\ll 1$), $\tilde{h}=\lambda\bar{J}$, and take $J_{j}^{(\text{int})}=0.1\bar{J}$, $\bar{J}=3\pi/8$ for concreteness. The $\pi$-pairing is exact for $\lambda=0$ and is expected to receive corrections $\sim\lambda^{L}$ for small $\lambda$ because it requires at least order $L$ in perturbation theory for the $W^{(e)}_{j,j+1}$ to combine to $\smash{\overline{W}}^{(e)}$, thus contributing to a splitting between PBC and APBC, while $\tilde{h}\tilde{X}_{\tilde{1}}$, by the $\pi$ pairing not being a degeneracy, can contribute only in even orders of $\tilde{h}$ and only to shifts of the $\pi$ pair but not their splitting. [We set $h_{j}^{(\text{int})}=0$ for simplicity; with $h_{j}^{(\text{int})}\sim\lambda$ the corrections to exact $\pi$ pairing would be $\sim\lambda^{L/2}$.] In contrast, $\tilde{h}\tilde{X}_{\tilde{1}}$ splits the static PM’s degeneracy between PBC and APBC sectors as $\sim\lambda$. Unlike this, the $\pi$ pairing in the $0\pi$PM receives corrections exponentially small in $L$, hence the absolute stability against $\tilde{Z}_{2}$ symmetry breaking. (One can more formally prove absolute stability in the $0\pi$PM using local unitary reasoning similar to that in Ref. von Keyserlingk et al., 2016.)
@@
We next test absolute stability numerically. We define von Keyserlingk et al. (2016) $\Delta_{0}^{i}=\varepsilon_{i+1}-\varepsilon_{i}$, with $\varepsilon_{i}$ the $i^{\text{th}}$ quasienergy of $U_{\text{F}}^{(e,\langle m\rangle)}$ and $\Delta_{\pi}^{i}=|\varepsilon_{i+M/2}-\varepsilon_{i}-\pi|$, where $M=2^{L+1}$ (the dimension of ¯$\overline{\mathcal{H}}_{\text{dyn}}$ including both PBC and APBC sectors), i.e., $i+M/2$ is halfway across the spectrum from $i$. We expect $\overline{\Delta}_{0}\sim e^{-sL}$ (with $s\sim\ln 2$ independent of $\lambda$) while $\overline{\Delta}_{\pi}\sim\lambda^{L}$; we focus on $\overline{\Delta}_{\pi}<\overline{\Delta}_{0}$ to be able to identify $\pi$ spectral pairs in the many-body spectrum. Our numerical results are shown in Fig. B1.
@@
We focused on breaking ~$\tilde{\mathbb{Z}}_{2}$ symmetry on a single site for simplicity: this required only one extra qubit, for $\tilde{Z}_{\tilde{1}}$, $\tilde{X}_{\tilde{1}}$). With just one gauged link, we could not demonstrate the robustness against $\mathbb{Z}_{2}$ symmetry breaking since with a $\mathbb{Z}_{2}$-breaking perturbation $H_{z}=\sum_{j}h_{j}^{z}Z_{j}$ and $h_{j}^{z}\sim\lambda$, the $\pi$ pairing would receive corrections already at order $\lambda^{3}$ due to $Z_{1}\tilde{Z}_{\tilde{1}}Z_{2}$, $Z_{1}$ and $Z_{2}$ being able to combine into $\smash{\overline{W}}^{(e)}=\tilde{Z}_{\tilde{1}}$ at order $\lambda^{3}$. However, upon gauging all links, $W^{(e)}_{j,j+1}\to Z_{j}\tilde{Z}_{\tilde{j}}Z_{j+1}$ and thus $\smash{\overline{W}}^{(e)}=\prod_{\tilde{j}=1}^{L}\tilde{Z}_{\tilde{j}}$, we would find that the corrections to $\pi$ pairing scale as $\lambda^{L}$ even with $H_{Z}$ present. Testing this would require doubling not the Hilbert space, but the number of qubits, and is thus numerically much more challenging.
@@ C.1 Duality twisted local Ising Hamiltonians and exact zero modes
and the nonlocal generators are $X_{l},\ Z_{l-1}Z_{l}$. We also include the link operator $\tilde{Z}_{l-1}$, labeling BCs, updating $Z_{l-1}Z_{l}\to Z_{l-1}\tilde{Z}_{\tilde{l}-1}Z_{l}$, hence $Z_{l-1}Y_{l}\to Z_{l-1}\tilde{Z}_{\tilde{l}-1}Y_{l}$. In what follows, we set $l=1$, without loss of generality, thus $Z_{l-1}\tilde{Z}_{\tilde{l}-1}Y_{l}=Z_{L}\tilde{Z}_{\tilde{L}}Y_{1}$ for an $L$-site system.
We next focus on non-interacting Ising systems, i.e., whose Hamiltonians involve terms from Eq. (4), but not their products. In this case, it is useful to introduce $2L$ Majorana operators $c_{i}$, and an auxiliary pair $\Gamma_{1,2}$ for $\tilde{Z}_{\tilde{L}}$, via JW transformation. We have
$X_{j}=ic_{2j-1}c_{2j}\ \forall j,\ Z_{j}Z_{j+1}=ic_{2j}c_{2j+1}\ (j\neq L),\\ Z_{1}Z_{L}=\prod_{j=1}^{L-1}Z_{j}Z_{j+1}=\prod_{j=1}^{L-1}ic_{2j}c_{2j+1},\ \tilde{Z}_{\tilde{L}}=i\Gamma_{1}\Gamma_{2},$ (5)
@@
$P\equiv\prod_{j=1}^{L}X_{j}=\prod_{j=1}^{L}(ic_{2j-1}c_{2j})\ \Rightarrow\ Z_{L}Z_{1}=-iPc_{2L}c_{1},\\ \!\!\!Z_{L}Y_{1}=P(ic_{2L}c_{2}),\ Z_{L}\tilde{Z}_{\tilde{L}}Y_{1}=P(i\Gamma_{1}\Gamma_{2})(ic_{2L}c_{2}).$ (6)
@@
Since all local $\mathbb{Z}_{2}$-symmetric (and BC preserving) terms commute with $P$ and $\tilde{Z}_{\tilde{L}}$, we can consider each of the, $P\tilde{Z}_{\tilde{L}}=p\tilde{z}\equiv s=\pm 1$ sectors separately (here $p,\tilde{z}$ are the eigenvalues; $s$ is their product); within these, a Hamiltonian $H_{s}$ that is a linear combination of the local terms [Eq. (4)] is quadratic in Majorana operators $c_{j}$. Now, since $s$ is a number, $c_{1}$ is absent from $H_{s}$: $\gamma_{1}\equiv c_{1}$ is an exact zero mode of $H_{s}$ (but not of $H$). Since $H_{s}$ is a quadratic fermion Hamiltonian, if it has one exact zero mode, it must have another. This is indeed the case: from
@@
where $S=P\tilde{Z}_{\tilde{L}}$. This is not fermion linear, not local, and not even Hermitian. However, the combination $\Psi_{0}=i\gamma_{2}\gamma_{1}$ has the following familiar features: it is Hermitian, $𝟙\Psi_{0}^{2}=\openone$ and it exactly commutes with $H$.
Noting that $\Gamma_{1}$ and $\gamma_{1}=c_{1}$ are each absent from $H_{s}$ and anticommute with $c_{j\neq 1}$ and with $Z_{L}Y_{1}\tilde{Z}_{\tilde{L}}$, we can replace $\gamma_{1}\to\Gamma_{1}$ in $\Psi_{0}$ to find that $\tilde{\gamma}_{2}=i\gamma_{2}\Gamma_{1}$ is also Hermitian, exactly commutes with $H$, and satisfies $𝟙\tilde{\gamma}_{2}^{2}=\openone$, $\{\tilde{\gamma}_{1},\tilde{\gamma}_{2}\}=0$. Recalling $\tilde{\gamma}_{1}=i\gamma_{1}\Gamma_{1}$, we have $\Psi_{0}=i\tilde{\gamma}_{2}\tilde{\gamma}_{1}$. Thus, in $\Psi_{0}$ we recognize the nonlocal integral of motion $\mathcal{W}_{\sigma}^{(a)}$ described in the main text.

### 28 Feb 2024 v2

@@ Appendix A Hilbert space, charges, boundary conditions, degeneracies
In the main text we have discussed how the Hilbert space $\mathcal{H}_{\text{dyn}}$ of the 1D system is embedded into the 2D topological order (TO) as the subspace with no anyons away from the boundary $B_{\text{dyn}}$. Here we discuss how $\mathcal{H}_{\text{dyn}}$ decomposes into symmetry sectors and is enlarged with twisted sectors. We also provide further discussion on twisted boundary conditions (BCs) and spectral degeneracies. Our approach bridges between the condensed-matter oriented Refs. Lichtman et al., 2021; Moradi et al., 2023, with Ref. Moradi et al., 2023 working on the infinite cylinder and considering twisted BCs, while Ref. Lichtman et al., 2021 working on a finite cylinder (thus having SymTFT access to certain spectral degeneracies) but without twisted BCs. Ref. Gaiotto and Kulp, 2021 establishes results analogous to those discussed in this Appendix from a high-energy theory perspective. See also Refs. Chatterjee and Wen, 2023a; Ji and Wen, 2020; Bhardwaj and Schafer-Nameki, ; Bhardwaj et al., ; Bhardwaj et al., .
@@
For finite Abelian $G$, we have $\mathcal{H}_{\text{dyn}}=\mathcal{H}_{1}^{\otimes L}$, with $\mathcal{H}_{1}=\text{span}(\{\ket{g},g\in G\})$. $\mathcal{H}_{\text{dyn}}$ splits into symmetry sectors as $\mathcal{H}_{\text{dyn}}=\oplus_{\alpha}\mathcal{H}_{\text{dyn}}^{(1,\alpha)}$, where subspaces $\mathcal{H}_{\text{dyn}}^{(1,\alpha)}$ are of states “charged” as $\alpha\in\text{Rep}\,G$ under the $G$ symmetry, $\smash{\overline{W}}^{(g,1)}\ket{\psi}=\chi_{\alpha}(g)\ket{\psi}$, |$\forall\ket{\psi}\in\mathcal{H}_{\text{dyn}}^{(1,\alpha)}$. Starting with the $G$-symmetric subspace $\mathcal{H}_{\text{dyn}}^{(1,1)}$, with $\smash{\overline{W}}^{(g,1)}\ket{\psi}=\ket{\psi}$, $\forall g\in G,\forall\ket{\psi}\in\mathcal{H}_{\text{dyn}}^{(1,1)}$, by $\smash{\overline{W}}^{(g,1)}W_{k}^{(1,\alpha)}=\chi_{\alpha}(g)W_{k}^{(1,\alpha)}\smash{\overline{W}}^{(g,1)}$, we have $\mathcal{H}_{\text{dyn}}^{(1,\alpha)}=W_{k}^{(1,\alpha)}\mathcal{H}_{\text{dyn}}^{(1,1)}$ (regardless of the choice of endpoint $k$). The operators $W_{k}^{(1,\alpha)}$ thus introduce charge $(1,\alpha)$ to $B_{\text{dyn}}$ and thus to the 1D system.
@@
The space $\mathcal{H}_{\text{dyn}}^{(g,\alpha)}=W_{j}^{(g,1)}\mathcal{H}_{\text{dyn}}^{(1,\alpha)}=W_{j}^{(g,1)}W_{k}^{(1,\alpha)}\mathcal{H}_{\text{dyn}}^{(1,1)}$ is of states obeying $g$-twisted BCs. In the main text we interpreted this via the “$G$-Aharonov Bohm” effect; here we provide a more direct interpretation: Consider $\mathcal{T}_{L}$, the translation of $B_{\text{dyn}}$, and with it $j$ and $k$, by $L$ sites, keeping the rest of the cylinder fixed. This wraps $W_{j}^{(g,1)}W_{k}^{(1,\alpha)}\equiv W^{(a)}$ [with $a=(g,\alpha)$] around the cylinder and, since reconnecting the resulting $a$ string (more precisely, ribbon) amounts to a braid, we have $\mathcal{T}_{L}W^{(a)}\mathcal{T}_{L}^{\dagger}=e^{i\theta_{aa}}W^{(a)}\smash{\overline{W}}^{(a)}$. Thus, by $\smash{\overline{W}}^{(a)}\ket{\psi}=\mathcal{T}_{L}^{\dagger}\ket{\psi}=\ket{\psi}$ for |$\ket{\psi}\in\mathcal{H}_{\text{dyn}}^{(1,1)}$, we find that $\mathcal{T}_{L}\ket{\psi}=\chi_{\alpha}(g)\ket{\psi}$ for |$\ket{\psi}\in\mathcal{H}_{\text{dyn}}^{(g,\alpha)}$: the BCs are twisted. Note that the link $\tilde{j}=\langle j,j\!+\!1\rangle$ where we insert the BC twist is a gauge choice: it can be moved to $\tilde{j}^{\prime}$ upon conjugating the Hamiltonian (or Floquet unitary) by $W_{jj^{\prime}}^{(g,1)}$. To consider all twisted sectors together, we may work in the enlarged Hilbert space ¯$\overline{\mathcal{H}}_{\text{dyn}}=\oplus_{\alpha,g}\mathcal{H}_{\text{dyn}}^{(g,\alpha)}$. When different BCs are coupled, as is the case in the presence of $\tilde{G}$-breaking perturbations, we must work with ¯$\overline{\mathcal{H}}_{\text{dyn}}$.
@@
Turning to degeneracies, first recall that a gapped, or more generally many-body localized (MBL), 1D system is characterized by a Lagrangian subgroup $\mathcal{A}$; the $W_{k,k+1}^{(a)}$ with $a\in\mathcal{A}$ form a maximal set of local, mutually commuting $G$-symmetric operators, defining the zero-correlation-length (or fixed-point, FP) limit of the gapped, or MBL, phase. The topological ground-state degeneracy on the cylinder, equal to $|\mathcal{A}\cap\mathcal{A}_{\text{ref}}|$, i.e., the number of distinct strings that can connect $B_{\text{ref}}$ and $B_{\text{dyn}}$ without creating anyons, while avoiding crossing domain wall $D^{(\sigma)}_{\text{ref}}$, yields the ground-state (or spectral, for MBL) degeneracy of the 1D system within a twisted sector $\oplus_{\alpha}\mathcal{H}_{\text{dyn}}^{(g,\alpha)}$. Strings $W_{j}^{(a)}$ ($a\in\mathcal{A}$, $a\notin\mathcal{A}_{\text{ref}}$) that can similarly connect $B_{\text{ref}}$ and $B_{\text{dyn}}$ while (part of them) crossing $D^{(\sigma)}_{\text{ref}}$ yield degeneracies between twisted sectors. In the Floquet setting, both types of degeneracies may split into topological spectral multiplets (such as $\pi$ spectral pairing as discussed for $G=\mathbb{Z}_{2}$), with structure set by the drive’s $\smash{\overline{W}}^{(b)}$.
@@ Appendix B Absolute stability in the $0\pi$PM
To study this, we consider $U_{\text{F}}^{(e,\langle m\rangle)}=\smash{\overline{W}}^{(e)}e^{-iH_{1}}e^{-iH_{0}}$ with $H_{0}=\sum_{j}J_{j}X_{j}+J_{j}^{(\text{int})}X_{j}X_{j+1}+\tilde{h}\tilde{X}_{1}$ and $H_{1}=\sum_{j}h_{j}W_{j,j+1}^{(e)}+h_{j}^{(\text{int})}W_{j,j+2}^{(e)}$. With only $\tilde{X}_{1}$ as link operator in $H_{0}$, it is sufficient to just gauge its link, i.e. to take $W_{j,j+1}^{(e)}=Z_{j}Z_{j+1}$ for $j\neq 1$ and $W_{1,2}^{(e)}=Z_{1}\tilde{Z}_{1}Z_{2}$, and $W^{(e)}_{L}=\tilde{Z}_{1}$. Since $\smash{\overline{W}}^{(e)}\tilde{X}_{1}=-\tilde{X}_{1}\smash{\overline{W}}^{(e)}$, the $\tilde{h}\tilde{X}_{1}$ term breaks ~$\tilde{\mathbb{Z}}_{2}$ symmetry; it couples the system’s untwisted, i.e., periodic boundary condition (PBC), and $m$-twisted, i.e., anti-periodic boundary condition (APBC), sectors.
@@
We take $J_{j}$ distributed uniformly within $[\bar{J}/2,3\bar{J}/2]$ and $h_{j}$ within $[-\lambda\bar{J},\lambda\bar{J}]$ (with $\lambda\ll 1$), $\tilde{h}=\lambda\bar{J}$, and take $J_{j}^{(\text{int})}=0.1\bar{J}$, $\bar{J}=3\pi/8$ for concreteness. The $\pi$-pairing is exact for $\lambda=0$ and is expected to receive corrections $\sim\lambda^{L}$ for small $\lambda$ because it requires at least order $L$ in perturbation theory for the $W^{(e)}_{j,j+1}$ to combine to $\smash{\overline{W}}^{(e)}$, thus contributing to a splitting between PBC and APBC, while $\tilde{h}\tilde{X}_{1}$, by the $\pi$ pairing not being a degeneracy, can contribute only in even orders of $\tilde{h}$ and only to shifts of the $\pi$ pair but not their splitting. [We set $h_{j}^{(\text{int})}=0$ for simplicity; with $h_{j}^{(\text{int})}\sim\lambda$ the corrections to exact $\pi$ pairing would be $\sim\lambda^{L/2}$.] In contrast, the static PM’s degeneracy between PBC and APBC sectors is split by $\tilde{h}\tilde{X}_{1}$ as $\sim\lambda$. Unlike this, the $\pi$ pairing in the $0\pi$PM receives corrections exponentially small in $L$, hence the absolute stability against $\tilde{Z}_{2}$ symmetry breaking. (One can more formally prove absolute stability in the $0\pi$PM using local unitary reasoning similar to that in Ref. von Keyserlingk et al., 2016.)
@@
We next test absolute stability numerically. We define von Keyserlingk et al. (2016) $\Delta_{0}^{i}=\varepsilon_{i+1}-\varepsilon_{i}$, with $\varepsilon_{i}$ the $i^{\text{th}}$ quasienergy of $U_{\text{F}}^{(e,\langle m\rangle)}$ and $\Delta_{\pi}^{i}=|\varepsilon_{i+M/2}-\varepsilon_{i}-\pi|$, where $M=2^{L+1}$ (the dimension of ¯$\overline{\mathcal{H}}_{\text{dyn}}$ including both PBC and APBC sectors), i.e., $i+M/2$ is halfway across the spectrum from $i$. We expect the typical values $\langle\ldots\rangle_{\text{typ}}=\exp(\langle\log(\ldots)\rangle)$ to satisfy $\langle\Delta_{0}\rangle_{\text{typ}}\sim e^{-sL}$ (with $s\sim\ln 2$ independent of $\lambda$) while $\langle\Delta_{\pi}\rangle_{\text{typ}}\sim\lambda^{L}$; we focus on $\Delta_{\pi}<\Delta_{0}$ to be able to identify $\pi$ spectral pairs in the many-body spectrum. Our numerical results are shown in Fig. B1.
@@
We focused on breaking ~$\tilde{\mathbb{Z}}_{2}$ symmetry on a single site for simplicity: this required only one extra qubit (for $\tilde{Z}_{1}$, $\tilde{X}_{1}$). With just one gauged link, we could not demonstrate the robustness against $\mathbb{Z}_{2}$ symmetry breaking since with a $\mathbb{Z}_{2}$-breaking perturbation $H_{z}=\sum_{j}h_{j}^{z}Z_{j}$ and $h_{j}^{z}\sim\lambda$, the $\pi$ pairing would receive corrections already at order $\lambda^{3}$ due to $Z_{1}\tilde{Z}_{1}Z_{2}$, $Z_{1}$ and $Z_{2}$ being able to combine into $\smash{\overline{W}}^{(e)}=\tilde{Z}_{1}$. However, upon gauging all links, $W^{(e)}_{j,j+1}\to Z_{j}\tilde{Z}_{j}Z_{j+1}$ and thus $\smash{\overline{W}}^{(e)}=\prod_{j=1}^{L}\tilde{Z}_{j}$, we would find that the corrections to $\pi$ pairing scale as $\lambda^{L}$ even with $H_{Z}$ present. Testing this would require doubling not the Hilbert space, but the number of qubits, and is thus numerically much more challenging.
@@ C.1 Duality twisted local Ising Hamiltonians and exact zero modes
and the nonlocal generators are $X_{l},\ Z_{l-1}Z_{l}$. We also include the link operator $\tilde{Z}_{l-1}$, labeling BCs, updating $Z_{l-1}Z_{l}\to Z_{l-1}\tilde{Z}_{l-1}Z_{l}$, hence $Z_{l-1}Y_{l}\to Z_{l-1}\tilde{Z}_{l-1}Y_{l}$. In what follows, we set $l=1$, without loss of generality, thus $Z_{l-1}\tilde{Z}_{l-1}Y_{l}=Z_{L}\tilde{Z}_{L}Y_{1}$ for an $L$-site system.
We next focus on non-interacting Ising systems, i.e., whose Hamiltonians involve terms from Eq. (4), but not their products. In this case, it is useful to introduce $2L$ Majorana operators $c_{i}$, and an auxiliary pair $\Gamma_{1,2}$ for $\tilde{Z}_{L}$, via JW transformation. We have
$X_{j}=ic_{2j-1}c_{2j}\ \forall j,\ Z_{j}Z_{j+1}=ic_{2j}c_{2j+1}\ (j\neq L),\\ Z_{1}Z_{L}=\prod_{j=1}^{L-1}Z_{j}Z_{j+1}=\prod_{j=1}^{L-1}ic_{2j}c_{2j+1},\ \tilde{Z}_{L}=i\Gamma_{1}\Gamma_{2},$ (5)
@@
$P\equiv\prod_{j=1}^{L}X_{j}=\prod_{j=1}^{L}(ic_{2j-1}c_{2j})\ \Rightarrow\ Z_{L}Z_{1}=-iPc_{2L}c_{1},\\ \!\!\!Z_{L}Y_{1}=P(ic_{2L}c_{2}),\ Z_{L}\tilde{Z}_{L}Y_{1}=P(i\Gamma_{1}\Gamma_{2})(ic_{2L}c_{2}).$ (6)
@@
Since all local $\mathbb{Z}_{2}$-symmetric (and BC preserving) terms commute with $P$ and $\tilde{Z}_{L}$, we can consider each of the, $P\tilde{Z}_{L}=p\tilde{z}\equiv s=\pm 1$ sectors separately (here $p,\tilde{z}$ are the eigenvalues; $s$ is their product); within these, a Hamiltonian $H_{s}$ that is a linear combination of the local terms [Eq. (4)] is quadratic in Majorana operators $c_{j}$. Now, since $s$ is a number, $c_{1}$ is absent from $H_{s}$: $\gamma_{1}\equiv c_{1}$ is an exact zero mode of $H_{s}$ (but not of $H$). Since $H_{s}$ is a quadratic fermion Hamiltonian, if it has one exact zero mode, it must have another. This is indeed the case: from
@@
where $S=P\tilde{Z}_{L}$. This is not fermion linear, not local, and not even Hermitian. However, the combination $\Psi_{0}=i\gamma_{2}\gamma_{1}$ has the following familiar features: it is Hermitian, $𝟙\Psi_{0}^{2}=\openone$ and it exactly commutes with $H$.
Noting that $\Gamma_{1}$ and $\gamma_{1}=c_{1}$ are each absent from $H_{s}$ and anticommute with $c_{j\neq 1}$ and with $Z_{L}Y_{1}\tilde{Z}_{L}$, we can replace $\gamma_{1}\to\Gamma_{1}$ in $\Psi_{0}$ to find that $\tilde{\gamma}_{2}=i\gamma_{2}\Gamma_{1}$ is also Hermitian, exactly commutes with $H$, and satisfies $𝟙\tilde{\gamma}_{2}^{2}=\openone$, $\{\tilde{\gamma}_{1},\tilde{\gamma}_{2}\}=0$. Recalling $\tilde{\gamma}_{1}=i\gamma_{1}\Gamma_{1}$, we have $\Psi_{0}=i\tilde{\gamma}_{2}\tilde{\gamma}_{1}$. Thus, in $\Psi_{0}$ we recognize the nonlocal integral of motion $\mathcal{W}_{\sigma}^{(a)}$ described in the main text.